40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 19 23:34:52 2012, kiwamu, Update, IOO, PMC realignment and HEPA 
    Reply  Mon Jan 23 15:43:47 2012, Jenne, Update, IOO, PMC realignment 
       Reply  Mon Jan 23 23:12:00 2012, kiwamu, Update, IOO, PMC realignment 
          Reply  Mon Jan 30 08:17:06 2012, steve, Update, IOO, PMC  sickPMC.png
             Reply  Tue Jan 31 08:44:10 2012, steve, Update, IOO, laser shuts down laseroff.png
                Reply  Tue Jan 31 14:58:30 2012, kiwamu, Update, IOO, laser shuts down 
                   Reply  Tue Jan 31 17:13:49 2012, steve, Update, IOO, only the PSL laser is off 
                      Reply  Wed Feb 1 17:00:57 2012, steve, Update, IOO, laser is back ON P1080525.JPGP1080514.JPGP1080518.JPG
                         Reply  Fri Feb 3 14:47:51 2012, steve, Update, PEM, laser interlock drawing 
Message ID: 6217     Entry time: Mon Jan 23 15:43:47 2012     In reply to: 6213     Reply to this: 6218
Author: Jenne 
Type: Update 
Category: IOO 
Subject: PMC realignment 

Quote:

I realigned the incident beam to PMC at 23:30. The transmitted light went up from 0.78 to 0.83.

 Do we have PSL pos and ang QPD trends?  We should start watching them, because the PMC drifted back down to 0.76 transmission, ~3.5 days after Kiwamu realigned it (his elog is from last Thurs).  Not so awesome.

I walked through the control room just now and found both PMC and MC unlocked.  They're both locked now, but with PMC transmission 0.76, MC transmission ~24,500.

ELOG V3.1.3-